Based on 6.4-13. Let Xi, . . . , Χη 1.1 a. Unif(9-1.0 + 2 (1) Find the method-of-moments estimator of θ (2) Is this estimatorSpecifically Part (3)
Answer #1

3)

The probability density function of X is given by:

f(0) = f(0) = 10 + 2) - (0 - 1) 1 1<x<0+2 3

The likelihood function is given by:

L = [[ f(x:) = which cannot be used to calculate MLE.

Now, 8-1< X (1) < X (2)... < X (n) <0+2 where X1, X2), ..., Xinh are the order statistics.

So, MLE is given by: A = mar(X (1) + 1, X(n)-2)

Thus, we can see that there are several possible choices for MLE for this problem.